At the end of the year, Mercy Cosmetics balance of Allowance for Uncollectible Accounts is $490


Brief Exercise 5-6 Record the adjusting for uncollectible accounts (L05-5) 


At the end of the year, Mercy Cosmetics balance of Allowance for Uncollectible Accounts is $490 (credid before adjustment. The balance of Accounts Receivable is $19,500. The company estimates that 10% of accounts will not be collected over the next year. 


What adjustment would Mercy Cosmetics record for Allowance for Uncollectible Accounts? (if no entry is required for a particular transaction/event, select "No Journal Entry Required" in the first account field.) 
Answer #1
1
General Journal Debit Credit
Bad Debts expense 1460 =(19500*10%)-490
       Allowance for Uncollectible accounts 1460
2
General Journal Debit Credit
Bad Debts expense 3154 =(26600*9%)+760
       Allowance for Uncollectible accounts 3154
3
Year Interest revenue
2021 500 =25000*8%*3/12
2022 1500 =25000*8%*9/12